Convert energy of 500 ergs into system based upon 100g,10cm, 30sec

Grade:11

2 Answers

Eshan
5 years ago
Letx=100g,y=10cm,z=30s
\implies 1g=\dfrac{x}{100},\implies 1cm=\dfrac{y}{10},\implies 1s=\dfrac{z}{30}

500 ergs=500\dfrac{g.cm^2}{s^2}=500\dfrac{\dfrac{x}{100}(\dfrac{y}{10})^2}{(\dfrac{z}{30})^2}
=45xy^2/z^2
Gitanjali Rout
5 years ago
500 ergs=500\dfrac{g.cm^2}{s^2}=500\dfrac{\dfrac{x}{100}(\dfrac{y}{10})^2}{(\dfrac{z}{30})^2}
hence hope that you find the answer useful
thankyou please  press upvite
thanks